OpenStudy (anonymous):

Solve 5m − 3h = 45 for h. Show your work.

OpenStudy (decentnabeel):

5m − 3h = 45 subtract 5m on both sides 5m-5m-3h=45-5m -3h=45-5m divided -3 on both sides -3h/-3=(45-5m)/-3 h=-45/3+5m/3 h=-15+5m/3

OpenStudy (decentnabeel):

are you understand @angie_doeee

OpenStudy (anonymous):

Yes.I understand.Im going to work it out the way you showed.

OpenStudy (decentnabeel):

ok

OpenStudy (anonymous):

hi